OpenStudy (anonymous):

Write an expression in simplest form for the volume of a cube with a side of 11x6 units

OpenStudy (goformit100):

@satellite73

OpenStudy (anonymous):

Well hello and welcome to openstudy! When you write 11x6 do you mean 11*6 (times) or 11x^6 (11 times x to the sixth power) ?

OpenStudy (anonymous):

1^6. My bad haha. And thank you!

OpenStudy (anonymous):

Sorry I'm still not sure what you mean, but I'll try to help. The volume of a cube is: \[V = a^{3}\] Where _a_ is the length of the side. If the side is 1^6 = 1 then the volume is 1^3 which is 1 too. If the side is 11*x^6 then the volume is (11*x^6)^3 = 1331*x^18 What ever the length of the side is, raise it to the power of 3 :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Your welcome =)